Help! Please Thanks!Cellphone Radiation Points:3 At a distance of 5 cm, the microwave radiation from a cellphone has an intensity of 29 W/m2. The maximum permissible radiation leaking out from a microwave oven at the same distance is 10 W/m2. Compare the average electric fields generated by the two devices - how many times stronger is the electric field that is generated by the cellphone than the maximum permissible field leaking from a microwave oven? 1.70 ou are correct. Your receipt no. is 164-3187 A bluetooth headset has an intensity of 0.08 W/m2 at S cm distance. If you carry your phone on your belt, 71 cm from the center of your brain, what is the total intensity (headset plus phone) at the center of your brain?0.000397 W/m 2 Submit Answer Incorrect. Tries 4/12 Previous Tries How many times longer than the width of your head (assume 10 cm) is the wavelength of a cellphone transmitting at 1753.2 MH2? 1.7 You are correct. Previous Tries ur receipt no. is 164-9921
Answer #1

I = intensity at 5 cm distance = 0.08 W/m2

r = 5 cm = 0.05 m

power radiated is given as

P = I A
P = I (4\pir2)

P = (0.08) (4 x 3.14 ) (0.05)2 = 0.002512 Watt

R = 71 cm = 0.71 m

Intensity of headset at the brain is given as

I = P/(4\piR2)

I = 0.002512/(4 x 3.14 (0.71)2)

I = 3.97 x 10-4 W/m2

for the phone :

I = 29 W/m2

r = 5 cm = 0.05 m

Power of phone is given as

P = I (4\pir2) = (29) (4 x 3.14) (0.05)2 = 0.9106 Watt

R = distance of brain = 71 cm = 0.71 m

I = intensity at the brain by phone = P/(4\piR2)

I = 0.9106/(4 x 3.14 (0.71)2) = 0.144 W/m2

Itotal = total intensity = 0.144 + 3.97 x 10-4 = 0.144 W/m2